# Cantor sets in higher dimensions II: Optimal dimension constraint for stable intersections

### v1

@@ Abstract.
It is well known that a pair of compact sets in $\mathbb{R}^{d}$ ($d\in\mathbb{N}$) can be separated by small deformations if the sum of their Hausdorff dimensions is less than $d$. In this paper, we demonstrate that this dimension constraint is optimal for regular Cantor sets. Specifically, for any prescribed Hausdorff dimensions whose sum is greater than $d$, we construct classes of pairs of regular Cantor sets that exhibit $C^{1+\alpha}$-stable intersections.
@@ 1. Introduction
$\operatorname{dim_{H}}K+\operatorname{dim_{H}}K^{\prime}<d,$
the set $K-K^{\prime}$ has zero Lebesgue measure. Consequently, almost all translations $K^{\prime}+t$ are disjoint from $K$.
The regime where $\operatorname{dim_{H}}K+\operatorname{dim_{H}}K^{\prime}>d$ is markedly different and considerably more subtle. Here, one expects not only positive measure but potentially the presence of interior points. This aligns with the Palis conjecture [13] regarding the arithmetic sum of regular Cantor sets on the real line. For such sets (including self-affine sets) one considers the stronger notion of stable intersection. This property implies that not only does $K-K^{\prime}$ possess a non-empty interior, but that this intersection persists robustly under small perturbations of the defining dynamical systems. Recall that a regular Cantor set, by definition, is the unique invariant set under the action of expanding smooth maps with restricted to a specific Markov partition. Two regular Cantor sets are $C^{r}$-close if their generating maps and Markov partitions are close in the $C^{r}$ topology and Hausdorff topology, respectively. For precise definitions we refer to [11].
@@
Let $d\in\mathbb{N}$ and $p,q\in(0,d)$ with $p+q>d$. Then there exists a pair of regular Cantor sets $(K,K^{\prime})$ in $\mathbb{R}^{d}$ with $\operatorname{dim_{H}}(K)=p$ and $\operatorname{dim_{H}}(K^{\prime})=q$ such that the pair $(K,K^{\prime})$ has $\mathcal{C}^{1+\alpha}$-stable intersection ($\alpha>0$).
@@
Figure 1. Dimension pairs $(\operatorname{dim_{H}}(K),\operatorname{dim_{H}}(K^{\prime}))$. Shaded region: pairs previously achieved by blender-type Cantor sets. Light gray region: new dimension pairs accessed via Theorems A and B.
@@ 4.1. Proof of Theorem A
By adding auxiliary affine generators to $K$ and choosing small and suitable $\delta,\varepsilon>0$, the intermediate value theorem allows us to arrange that $\dim_{\mathrm{Aff}}(K)=p$, using the continuity of the affinity dimension with respect to the affine generators [6]. Although $\dim_{\mathrm{Aff}}(K)$ need not coincide with $\dim_{\mathrm{H}}(K)$ in general, Falconer’s theorem [5] ensures that, after a small perturbation of the translation parts of the generators, equality holds. Moreover, the stable intersection property is preserved under this perturbation, since the covering criterion is open. The same idea of auxiliary generators also helps to construct $K^{\prime}$ such that one has $\dim_{\operatorname{\rm{Aff}}}(K^{\prime})=\operatorname{dim_{H}}(K^{\prime})=q>d-q+\delta$. ∎
@@ 4.2. Proof of Theorem B
Step 4. By Falconer’s theorem [5], for almost every choice of vectors $v_{i},v^{\prime}_{j^{\prime}}\in\mathbb{R}^{d}$, the affinity dimension and Hausdorff dimension coincide. Since stable intersection is an open property with respect to the Cantor sets $K$ and $K^{\prime}$, we can perturb the vectors $v_{i},v^{\prime}_{j^{\prime}}\in\mathbb{R}^{d}$ so that
$\dim_{\rm{H}}(K)=\dim_{\rm{Aff}}(K),\quad\dim_{\rm{H}}(K^{\prime})=\dim_{\rm{Aff}}(K^{\prime}).$
@@
$\dim_{\rm{Aff}}(K_{d})=\dim_{\rm{H}}(K_{d})=d\cdot\dim_{\rm{H}}(K_{1})\in\left(p+\frac{\varepsilon}{3},p+\frac{2\varepsilon}{3}\right),$
$\dim_{\rm{Aff}}(K^{\prime}_{d})=\dim_{\rm{H}}(K^{\prime}_{d})=d\cdot\dim_{\rm{H}}(K^{\prime}_{1})\in\left(d-p+\frac{\varepsilon}{3},d-p+\frac{2\varepsilon}{3}\right).$
@@
$\operatorname{dim_{H}}(K)=\dim_{\operatorname{\rm{Aff}}}(K)=p,\;\;\operatorname{dim_{H}}(K^{\prime})=\dim_{\operatorname{\rm{Aff}}}(K^{\prime})=q$
@@ 5. Homogeneous Cantor sets on the real line
  • There exist $a\in\mathbb{R}^{+}$ and an interval $I\subset\mathbb{R}$ such that the set $\{a\}\times I$ is a recurrent set under the action of renormalization operators of the pair $(K_{1},K_{1}^{\prime})$. Indeed, $\{a\}\times I$ satisfies $c$-separable covering condition with respect to the renormalization operators of the pair $(K_{1},K_{1}^{\prime})$.

### v2

@@ Abstract.
It is well known that a pair of compact sets in $\mathbb{R}^{d}$ ($d\in\mathbb{N}$) can be separated by small deformations if the sum of their upper box dimensions is less than $d$. In this paper, we demonstrate that this dimension constraint is optimal for regular Cantor sets. Specifically, for any prescribed upper box dimensions whose sum is greater than $d$, we construct classes of pairs of regular Cantor sets that exhibit $C^{1+\alpha}$-stable intersections.
@@ 1. Introduction
$\operatorname{\overline{dim}_{B}}(K)+\operatorname{\overline{dim}_{B}}(K^{\prime})<d,$
the set $K-K^{\prime}$ has zero Lebesgue measure. Consequently, almost all translations $K^{\prime}+t$ are disjoint from $K$. Here, $\operatorname{\overline{dim}_{B}}$ denotes the upeer box dimension (see §2).
The regime where $\operatorname{\overline{dim}_{B}}(K)+\operatorname{\overline{dim}_{B}}(K^{\prime})>d$ is markedly different and considerably more subtle. Here, one expects not only positive measure but potentially the presence of interior points. This aligns with the Palis conjecture [13] regarding the arithmetic sum of regular Cantor sets on the real line. For such sets (including self-affine sets) one considers the stronger notion of stable intersection. This property implies that not only does $K-K^{\prime}$ possess a non-empty interior, but that this intersection persists robustly under small perturbations of the defining dynamical systems. Recall that a regular Cantor set, by definition, is the unique invariant set under the action of expanding smooth maps with restricted to a specific Markov partition. Two regular Cantor sets are $C^{r}$-close if their generating maps and Markov partitions are close in the $C^{r}$ topology and Hausdorff topology, respectively. For precise definitions we refer to [11].
@@
Let $d\in\mathbb{N}$ and $p,q\in(0,d)$ with $p+q>d$. Then there exists a pair of regular Cantor sets $(K,K^{\prime})$ in $\mathbb{R}^{d}$ with $\operatorname{\overline{dim}_{B}}(K)=p$ and $\operatorname{\overline{dim}_{B}}(K^{\prime})=q$ such that the pair $(K,K^{\prime})$ has $\mathcal{C}^{1+\alpha}$-stable intersection ($\alpha>0$).
@@
Figure 1. Dimension pairs $(\operatorname{\overline{dim}_{B}}(K),\operatorname{\overline{dim}_{B}}(K^{\prime}))$. Shaded region: pairs previously achieved by blender-type Cantor sets. Light gray region: new dimension pairs accessed via Theorems A and B.
@@ 2. Preliminaries
  • We denote the upper box dimension of the set $A\subset\mathbb{R}^{d}$ as
    $\operatorname{\overline{dim}_{B}}(A):=\limsup_{\varepsilon\to 0}\frac{\log N(A,\varepsilon)}{\log(1/\varepsilon)},$
    where $N(A,\varepsilon)$ is the minimum number of sets of diameter $\varepsilon$ required to cover $A$.
@@ 4.1. Proof of Theorem A
By adding auxiliary affine generators to $K$ and choosing small and suitable $\delta,\varepsilon>0$, the intermediate value theorem allows us to arrange that $\dim_{\mathrm{Aff}}(K)=p$, using the continuity of the affinity dimension with respect to the affine generators [6]. Although $\dim_{\mathrm{Aff}}(K)$ need not coincide with $\operatorname{\overline{dim}_{B}}(K)$ and $\dim_{\mathrm{H}}(K)$ in general, Falconer’s theorem [5] ensures that, after a small perturbation of the translation parts of the generators, equality holds. Moreover, the stable intersection property is preserved under this perturbation, since the covering criterion is open. The same idea of auxiliary generators also helps to construct $K^{\prime}$ such that one has $\operatorname{\overline{dim}_{B}}(K^{\prime})=\dim_{\operatorname{\rm{Aff}}}(K^{\prime})=\operatorname{dim_{H}}(K^{\prime})=q>d-q+\delta$. ∎
@@ 4.2. Proof of Theorem B
Step 4. By Falconer’s theorem [5], for almost every choice of vectors $v_{i},v^{\prime}_{j^{\prime}}\in\mathbb{R}^{d}$, the upper box dimension and Hausdorff dimension coincide with the affinity dimension. Since stable intersection is an open property with respect to the Cantor sets $K$ and $K^{\prime}$, we can perturb the vectors $v_{i},v^{\prime}_{j^{\prime}}\in\mathbb{R}^{d}$ so that
$\operatorname{dim_{H}}(K)=\operatorname{\overline{dim}_{B}}(K)=\operatorname{dim_{Aff}}(K),\quad\operatorname{dim_{H}}(K^{\prime})=\operatorname{\overline{dim}_{B}}(K^{\prime})=\operatorname{dim_{Aff}}(K^{\prime}).$
@@
$\operatorname{\overline{dim}_{B}}(K_{d})=\dim_{\rm{Aff}}(K_{d})=\dim_{\rm{H}}(K_{d})=d\cdot\dim_{\rm{H}}(K_{1})\in\left(p+\frac{\varepsilon}{3},p+\frac{2\varepsilon}{3}\right),$
$\operatorname{\overline{dim}_{B}}(K^{\prime}_{d})=\dim_{\rm{Aff}}(K^{\prime}_{d})=\dim_{\rm{H}}(K^{\prime}_{d})=d\cdot\dim_{\rm{H}}(K^{\prime}_{1})\in\left(d-p+\frac{\varepsilon}{3},d-p+\frac{2\varepsilon}{3}\right).$
@@
$\operatorname{\overline{dim}_{B}}(K)=\operatorname{dim_{H}}(K)=\dim_{\operatorname{\rm{Aff}}}(K)=p,\;\;\operatorname{\overline{dim}_{B}}(K^{\prime})=\operatorname{dim_{H}}(K^{\prime})=\dim_{\operatorname{\rm{Aff}}}(K^{\prime})=q$
@@ 5. Homogeneous Cantor sets on the real line
  • There exist $a\in\mathbb{R}^{+}$ and an interval $I\subset\mathbb{R}$ such that the set $\{a\}\times I$ is a recurrent set under the action of renormalization operators of the pair $(K_{1},K_{1}^{\prime})$. Indeed, $\{a\}\times I$ satisfies the $c$-separable covering condition with respect to the renormalization operators of the pair $(K_{1},K_{1}^{\prime})$.